A 0.045 0 M solution of HA is 0.60% dissociated. Calculate pKa for this acid.


1
Expert's answer
2021-06-18T07:33:03-0400

Solution:

Dissociation reaction: HA ⇄ H+ + A-

The Ka expression: Ka = [H+][A-] / [HA]


ICE Table:

___________HA___⇄__H+__+__A-__

Initial:_____0.045_____0______0___

Change:____-x_______+x_____+x___

Equlib:____0.045-x____x______x___


Ka = [H+][A-] / [HA]

Ka = (x)(x) / (0.045 - x)

0.045 - x ≈ 0.045 (assuming that 0.045 >> x)

Therefore,

Ka = x2 / 0.045


Since acid is 0.60% dissociated,

0.006 = [H+] / [HA] = x / 0.045

x = (0.006)(0.045) = 0.00027

x = 0.00027


Substitute 'x' in Ka = x2 / 0.045 to get Ka:

Ka = (0.00027)2 / 0.045 = 1.62×10-6

Ka = 1.62×10-6

pKa = -log(Ka) = -log(1.62×10-6) = 5.79

pKa = 5.79


Answer: pKa = 5.79